OpenStudy (anonymous):

Cylinder A has radius 1 m and height 4 m. Cylinder B has radius 2 m and height 4 m. Find the ratio of the volume of cylinder A to the volume of cylinder B.

OpenStudy (kropot72):

Volume of a cylinder = \[\ \pi r ^{2}h\] \[V _{a}=\ \pi \times 1^{2}\times 4\] \[V _{b}=\pi \times 2^{2}\times 4\] \[\frac{V _{a}}{V _{b}}=\frac{\ \pi \times 4}{\pi \times 4\times 4}=?\]

OpenStudy (anonymous):

thank you!!!

OpenStudy (kropot72):

You're welcome :)